# Zero-shot LLM-guided Counterfactual Generation for Text

### v1

@@ Abstract
Counterfactual examples are frequently used for model development and evaluation in many natural language processing (NLP) tasks. Although methods for automated counterfactual generation have been explored, such methods depend on models such as pre-trained language models that are then fine-tuned on auxiliary, often task-specific datasets. Collecting and annotating such datasets for counterfactual generation is labor intensive and therefore, infeasible in practice. Therefore, in this work, we focus on a novel problem setting: zero-shot counterfactual generation. To this end, we propose a structured way to utilize large language models (LLMs) as general purpose counterfactual example generators. We hypothesize that the instruction-following and textual understanding capabilities of recent LLMs can be effectively leveraged for generating high quality counterfactuals in a zero-shot manner, without requiring any training or fine-tuning. Through comprehensive experiments on various downstream tasks in natural language processing (NLP), we demonstrate the efficacy of LLMs as zero-shot counterfactual generators in evaluating and explaining black-box NLP models.
[…]
2725
1 Introduction
Over the last couple of decades, machine learning and natural language processing (NLP) systems have developed massively, especially in terms of the complexity and scale of the models used in different downstream tasks. For example, for most NLP tasks, such as tasks in the GLUE or SuperGLUE benchmark, the state-of-the-art performance is achieved by large, black-box models such as pre-trained language models (PLM) [23]. Effective use and deployment of such models, especially in high stakes areas, require careful evaluation, validation and stress-testing. Furthermore, models should also be explainable or interpretable, i.e., decisions made by such black-box models should ideally be accompanied by how and/or why the model reached that decision [28]. While such endeavors are still challenging in the context of black-box models, in this regard, counterfactual examples have been used to perform evaluation, explanation, robustness testing and even improvement of NLP models [48, 26, 5]. For example, the following two sentences - s1: This movie is brilliant!, s2: This movie is boring. are counterfactual examples for the input sentence This movie is great. Such minimally perturbed variations of the input text can be used in a variety of settings to evaluate models, to understand whether a model is able to focus on the task-specific features in the input text in order to classify the input text correctly.
@@ 1 Introduction
While several previous works have investigated the applicability of human expert annotators to design such counterfactual examples [16, 32], this is not scalable in practice, thereby motivating the exploration of automated counterfactual generation methods. Automated counterfactual generation methods such as  [48, 26] use pre-trained language models, or mask-filling, or models trained via control codes for the generation. Such models require training or fine-tuning using multiple datasets that are often task-specific in nature. For example, training a conditional generation model in Polyjuice [48] requires sentence-pair dataset for each control code (such as: negation, quantifier, shuffle, lexical, etc.). Similar methods requiring vast amounts of training and/or data are used in other automated counterfactual generation methods. However, having access to such task-specific training datasets may be infeasible in practice, especially for newly emerging data domains and tasks. Therefore, we are interested in investigating: Is there a way to simplify the counterfactual generation process and perform the generation without any auxiliary data?
@@
To this end, in this work, we address a new problem setting: zero-shot counterfactual generation. We tackle this problem by using the power of recent state-of-the-art instruction-tuned large language models (LLMs). Given that these LLMs are trained on massive amounts of text data, followed by subsequent supervised fine-tuning and alignment steps, there is empirical evidence to suggest that such LLMs can be used as pseudo-oracles or general purpose solvers especially in natural language processing tasks [8]. As an extension, we propose the paradigm of using LLMs as zero-shot counterfactual generators for text. To this end, we propose a pipeline that leverages recent LLMs in order to generate plausible, human-interpretable counterfactual examples in a completely zero-shot manner. Our proposed pipeline requires only the input text along with either the ground truth label or the predicted label from the black-box classifier (depending on the use-case) and uses a simple hard-prompting method to use off-the-shelf LLMs for generating the counterfactuals, without any fine-tuning or training with additional data. We envision that automating the task of counterfactual generation via a carefully designed pipeline that leverages LLMs can help to reduce costs and make NLP model development, evaluation and explanation more streamlined and efficient. We use our proposed pipeline to generate counterfactual explanations and we empirically evaluate the quality and effectiveness of the generated counterfactuals in (1) explaining and, (2) evaluating NLP models for a variety of downstream tasks. Our results demonstrate that, when used in our pipeline, LLMs can generate high quality and effective zero-shot counterfactuals that preserve high semantic similarity with the original input text.
@@
To the best of our knowledge, this is the first piece of work to tackle the problem of zero-shot counterfactual generation in text, and also the first to utilize LLMs as general purpose counterfactual generators. Overall our contributions in this paper are as follows:
@@
    We propose a Framework for Instructed Zero-shot Counterfactual Generation with LanguagE Models, which we refer to as FIZLE for brevity [^1] All code and data will be available at ¡link-to-be-inserted-after-blind-review¿..
@@
    We demonstrate FIZLE for two important use-cases: explaining and evaluating black-box text classification models.
@@
    Through experiments on three benchmark datasets and tasks we investigate the effectiveness of the proposed pipeline compared to recent baselines and discuss implications for future work in this direction.
Figure 2: Our proposed FIZLE framework for zero-shot LLM-guided general purpose counterfactual generation, along with applications in model evaluation and explanation.
The rest of the paper is organized as follows: Section 2 dives into the background and relevant related works. Section 3 describes our pipeline and methodology in detail. We briefly talk about our general experimental settings in Section 4. We demonstrate our two use-cases, Counterfactual Explanations in Section 5 and Evaluation using Counterfactuals as Contrast Sets in Section 6. We finally conclude with a discussion on future works in Section 7.
2 Background and Related Works
In this section we describe some preliminary concepts along with related works.
Counterfactual Generation
According to most definitions in literature [28], counterfactuals in text are minimally edited versions of an original text that can flip the label of a classifier. Counterfactuals are typically similar to the input instance, and vary from it in a small number of features. Counterfactual examples are widely used to stress-test trained models, provide explanations in the form of counterfactual explanations, and also for model improvement via training with counterfactual examples and counterfactually augmented data. While several efforts have been made in the manual creation of counterfactuals [16, 32], this method does not scale up and is therefore infeasible in practice for most use-cases. Automated methods for counterfactual generation are therefore more prevalent. Such methods often use language models trained on some control codes for conditional text generation in order to generate plausible and diverse counterfactuals [26, 48]. In the context of text classification, which is the scope of this paper, counterfactual examples can be generated from the input text via token-based substitution methods, masked-language modeling, controlled text generation via control codes [27]. Authors in  [37] create realistic counterfactuals via language modeling using a Counterfactual GAN architecture. However, all of these methods use either auxiliary models and/or training data, for example, to capture the style characteristics of different control codes. In contrast, in this work we focus on zero-shot counterfactual generation.
@@ 2 Background and Related Works
Large Language Models (LLMs) are usually transformer-based models capable of generating human-like text. Recent examples of large language models include the GPT family of models [33, 7, 29], Llama [41], Llama-2 [42], Falcon [^2] https://huggingface.co/tiiuae/falcon-40b, etc. A general training recipe for training LLMs include an unsupervised pre-training step, where the model is trained using a huge corpora of text, typically comprising of text from the internet [31, 15], followed by one or more supervised fine-tuning steps, such as instruction-tuning [30, 12]. More recent state-of-the-art LLMs such as ChatGPT-3.5 or GPT-4 are further fine-tuned via reinforcement learning with human feedback (RLHF) [11], in order to ‘align’ such models more with human preferences and values. During the fine-tuning and RLHF stages LLMs learn to follow instructions for specific tasks and respond in a helpful manner. Instruction-tuning essentially fine-tunes the model on massive datasets of (instruction, output) pairs, whereby LLMs learn to follow instructions in a prompt in order to perform tasks. The vast amount of training, both via the pre-training and the instruction-tuning stages, enables the LLMs to perform complex tasks [8], perform in-context learning [14], etc. Recent advancements in LLMs have sparked simultaneous exploration and research into the applicability of these LLMs on a variety of different tasks, such as data labeling  [19, 3, 40], text classification  [39, 4], model explanation [5], etc. We add on to this emerging body of work and perform zero-shot general purpose counterfactual generation using LLMs.
3 FIZLE: Zero-shot LLM-guided Counterfactual Generation
In this section, we describe our main framework as shown in Figure 2. Following the causal explanation generation procedure in prior work  [5], we use state of the art LLMs in an off-the-shelf manner, without any fine-tuning. We improve upon prior work  [5] by expanding and broadening their pipeline into a more general framework that can work for tasks other than causal explanation. Note that in this paper, we formulate and evaluate our pipeline on the broad task of text classification, whereas formulations for other text tasks can be derived in a similar manner. To facilitate this we explain the following components in our framework:
@@ 3 FIZLE: Zero-shot LLM-guided Counterfactual Generation
We leverage recent state-of-the-art LLMs as the counterfactual generators. Given that these models have been trained on vast amounts of textual data along with extensive instruction tuning, we assume that LLMs can learn to modify and perturb text input to simulate how human annotators generate counterfactuals for specific tasks  [18]. For this, we use both proprietary models from OpenAI and open-source models from Meta AI, and use carefully crafted instructions and constraint prompts to generate the counterfactuals. Specifically, we use the following proprietary models via the OpenAI API where applicable:
  • text-davinci-003: This model was a pre-cursor to OpenAI’s GPT-3.5.
  • GPT-3.5333https://platform.openai.com/docs/models/gpt-3-5-turbo: Often referred to as ChatGPT. This is the model that has been explored in a variety of text applications. Specifically we use the gpt-3.5-turbo version as of January-February 2024.
@@
Among the open-source models, we use two sizes of the Llama 2 model wherever applicable. These are:
  • Llama 2 7B: This is the 7 billion parameter version of the model. We specifically use the ‘chat’ variant via Huggingface [^5] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.
  • Llama 2 13B: This is the 13 billion parameter version of the model. Similar to the previous one, we use the ‘chat’ variant from Huggingface [^6] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf.
Due to resource constraints, we were unable to use the largest 70B Llama 2 model.
@@
To generate the counterfactuals in a zero-shot manner using the chosen LLM, the prompt needs to have informative instructions and constraints to guide the generation. We generate two types of counterfactuals: (1) counterfactuals with different label from original: these are used in the counterfactual explanation experiments (see Section 5), and (2) counterfactuals with same label as original: these are used in the contrast set experiments (see Section 6). For setting (1), we experiment with two variants of the generation process: (i) naive: Here the LLM is directly prompted to generate a counterfactual, and (ii) guided: Here we use a two-step process - first leveraging the LLM to identify the important input features (i.e., words) that result in the predicted label, and then prompting the same LLM to edit a minimal set of those identified features to generate the counterfactual. We do this since generating a counterfactual that flips the label of the original input is intuitively more challenging and providing additional guidance to the LLM in the form of steps may potentially improve the quality of the generated counterfactuals. For ease of extraction of the generated rationales, we also specify an output constraint that allows easy parsing based on a regular expression string match.
4 Experimental Settings
We use our proposed FIZLE pipeline to generate counterfactuals and demonstrate these in two specific use-cases: (1) Counterfactual explanations for explaining decisions of black-box text classifiers and (2) Evaluating black-box text classification models via contrast sets. To facilitate this, here we go over the datasets used and the general experimental setup for all our generation and evaluation experiments:
4.1 Datasets
@@
4.2 Experimental Setup
All experiments on open-source models were performed on two A100 GPUs with a total of 80G memory. For 13B Llama 2 models, we use 4-bit quantization using the optimal ‘nf4’ datatype [13]. For all LLMs, we use top_p sampling with $p=1$, temperature $t=0.4$ and a repetition penalty of $1.1$. We use PyTorch and fine-tuned models hosted on Huggingface for in both Sections 5 and 6.
5 Counterfactual Explanations via LLM-generated Counterfactuals
Explainability is a major challenge in many NLP applications such as text classification [35, 2, 9]. Although recent models involving pre-trained transformer based language models [44] have achieved or even exceeded human-level performance on several tasks [45, 46], most of these models are black-box by design and hence are not interpretable. Such models do not offer transparency on why it predicted a certain label, or even what features in the input resulted in the prediction. The ubiquity of these black-box classifiers necessitates the development of explanation frameworks and techniques that provide some degree of understanding into the decision-making function of the model [1]. Counterfactual explanations [28] give an insight into what could have been different in the input to change the output label predicted by the classifier. Gold-standard counterfactual generation requires human annotators and is also task-specific [20, 21], therefore making it an extremely expensive and labor-intensive endeavor. Therefore, we use our LLM-genearted counterfactuals in the of counterfactual explanations for black-box text classifiers.

### v2

# Zero-shot LLM-guided Counterfactual Generation: A Case Study on NLP Model Evaluation

@@ Abstract
With the development and proliferation of large, complex, black-box models for solving many natural language processing (NLP) tasks, there is also an increasing necessity of methods to stress-test these models and provide some degree of interpretability or explainability. While counterfactual examples are useful in this regard, automated generation of counterfactuals is a data and resource intensive process. such methods depend on models such as pre-trained language models that are then fine-tuned on auxiliary, often task-specific datasets, that may be infeasible to build in practice, especially for new tasks and data domains. Therefore, in this work we explore the possibility of leveraging large language models (LLMs) for zero-shot counterfactual generation in order to stress-test NLP models. We propose a structured pipeline to facilitate this generation, and we hypothesize that the instruction-following and textual understanding capabilities of recent LLMs can be effectively leveraged for generating high quality counterfactuals in a zero-shot manner, without requiring any training or fine-tuning. Through comprehensive experiments on a variety of propreitary and open-source LLMs, along with various downstream tasks in NLP, we explore the efficacy of LLMs as zero-shot counterfactual generators in evaluating and explaining black-box NLP models.
Index Terms:
counterfactual generation, model evaluation, explanation, explainability, large language models
I Introduction
Over the last couple of decades, machine learning and natural language processing (NLP) systems have developed massively, especially in terms of the complexity and scale of the models used in different downstream tasks. For example, for most NLP tasks, such as tasks in the GLUE [1] or SuperGLUE [2] benchmarks, the state-of-the-art performance is achieved by large, black-box models such as pre-trained language models (PLM) [3]. Effective use and deployment of such models, especially in high-stakes areas, require careful evaluation, validation and stress-testing. Furthermore, models should also be explainable or interpretable, i.e., decisions made by such black-box models should ideally be accompanied by how and/or why the model reached that decision [4]. While such endeavors are still challenging in the context of black-box models, in this regard, counterfactual examples have been used to perform evaluation, explanation, robustness testing and even improvement of NLP models [5, 6, 7]. For example, the following two sentences - s1: This movie is brilliant!, s2: This movie is boring. are counterfactual examples for the input sentence This movie is great. Such minimally perturbed variations of the input text can be used in a variety of settings to evaluate models, to understand whether a model is able to focus on the task-specific features in the input text in order to classify the input text correctly.
@@ I Introduction
While several previous works have investigated the applicability of human expert annotators to design such counterfactual examples [8, 9], this is not scalable in practice, thereby motivating the exploration of automated counterfactual generation methods. Automated counterfactual generation methods such as  [5, 6] use pre-trained language models, or mask-filling, or models trained via control codes for the generation. For example, training a conditional generation model in Polyjuice [5] requires sentence-pair dataset for each control code (such as: negation, quantifier, shuffle, lexical, etc.). Similar methods requiring large amounts of training and/or task-specific data are used in other automated counterfactual generation methods. However, having access to such task-specific training datasets may be infeasible in practice, especially for newly emerging data domains and tasks. Therefore, we are interested in investigating: Is there a way to simplify the counterfactual generation process and perform the generation without any auxiliary data?
@@
In order to explore this, in this work, we address a new problem and application setting: zero-shot counterfactual generation for evaluating and explaining NLP models. We tackle this problem by using the power of recent state-of-the-art instruction-tuned large language models (LLMs). While recent work has started exploring the effectiveness of using LLMs for generating counterfactuals, these works use additional guidance, such as, in the form of few-shot exemplars for in-context learning. Given that gold-standard exemplar samples can be hard to come across for many tasks, alongside LLM context length issues restricting the number of in-context examples, we explore the possibility of using LLMs in a zero-shot manner for generating counterfactuals in order to evaluate and explain black-box text classifiers. Given that recent LLMs are trained on massive amounts of text data, followed by subsequent supervised fine-tuning and alignment steps, empirical evidence suggests that such LLMs can be used as pseudo-oracles or general-purpose solvers especially in NLP tasks [10]. As an extension, we propose the paradigm of using LLMs as zero-shot counterfactual generators for stress-testing text classifier models. To further this exploration, we propose a pipeline that leverages recent LLMs in order to generate plausible, human-interpretable counterfactual examples in a completely zero-shot manner. Our proposed pipeline requires only the input text along with either the ground truth label or the predicted label from the black-box classifier (depending on the use-case) and uses a structured, hard-prompting method to use off-the-shelf LLMs for generating the counterfactuals, without any fine-tuning or training with additional data. We envision that automating the task of counterfactual generation via a carefully designed pipeline that leverages LLMs can help to reduce costs and make NLP model development, evaluation and explanation more streamlined and efficient. We use our proposed pipeline to generate counterfactuals in order to explore their effectiveness in (1) explaining and, (2) evaluating NLP models for a variety of downstream tasks. Our results demonstrate that, when used in our pipeline, LLMs may be effectively used to generate effective zero-shot counterfactuals that can be used for stress-testing text classifiers.
@@
To the best of our knowledge, this is the first piece of work to tackle the problem of zero-shot counterfactual generation to evaluate and explain text classifiers. Overall our contributions in this paper are as follows:
@@
    We explore a Framework for Instructed Zero-shot Counterfactual Generation with LanguagE Models, which we refer to as FIZLE for brevity  [^1] All code, prompts, supplementary materials, etc. are available at https://github.com/AmritaBh/zero-shot-llm-counterfactual..
@@
    We investigate and evaluate FIZLE for two important use-cases: explaining and evaluating black-box text classification models.
@@
    Through experiments on three benchmark datasets, several open-source and proprietary LLMs, we investigate the effectiveness of the proposed pipeline compared to recent baselines and discuss implications for future work in this direction.
Figure 2: Our proposed FIZLE pipeline for zero-shot LLM-guided counterfactual generation for evaluation and explanation of black-box text classifiers.
II Background and Related Works
In this section we describe some preliminary concepts along with relevant related works.
Counterfactual Examples
According to most definitions in literature [4], counterfactuals in text are minimally edited versions of an original text that can flip the label of a classifier222although, in this paper, we often use the phrase “counterfactual with same label as input” which effectively refers to a semantically similar re-write of the input text, i.e, similar to having undergone a label-preserving data augmentation step.. Counterfactuals are typically similar to the input instance, and vary from it in a small number of features. Counterfactual examples may be used to stress-test trained models, provide explanations in the form of counterfactual explanations, and also for model improvement via training with counterfactual examples and counterfactually augmented data. While several efforts have been made in the manual creation of counterfactuals [8, 9], this method does not scale up and is therefore infeasible in practice for most use-cases. Automated methods for counterfactual generation are therefore more prevalent. Such methods often use language models trained on some control codes for conditional text generation in order to generate plausible and diverse counterfactuals [6, 5]. In the context of text classification, which is the scope of this paper, counterfactual examples can be generated from the input text via token-based substitution methods, masked-language modeling, controlled text generation via control codes [11]. Authors in  [12] create realistic counterfactuals via language modeling using a Counterfactual GAN architecture. However, all of these methods use either auxiliary models and/or training data, for example, to capture the style characteristics of different control codes. Some recent work has also explored using large language models in few-shot settings for counterfactual generation [13]. Unlike previous work, in this paper, we focus on zero-shot counterfactual generation for stress-testing text classifiers.
@@ II Background and Related Works
Large Language Models (LLMs) are usually transformer-based models capable of generating human-like text. Recent examples of large language models include the GPT family of models from OpenAI [14, 15, 16], the Llama family of models from Meta [17, 18, 19], Gemini from Google  [20], etc. A general training recipe for training LLMs include an unsupervised pre-training step, where the model is trained using a huge corpora of text, typically comprising of text from the internet [21, 22], followed by one or more supervised fine-tuning steps, such as instruction-tuning [23, 24]. More recent state-of-the-art LLMs such as GPT-3.5 or GPT-4 are further fine-tuned via reinforcement learning with human feedback (RLHF) [25], in order to ‘align’ such models more with human preferences and values. During the fine-tuning and RLHF stages LLMs learn to follow instructions for specific tasks and respond in a helpful manner. Instruction-tuning essentially fine-tunes the model on massive datasets of (instruction, response) pairs, whereby LLMs learn to follow instructions in a prompt in order to perform tasks. The vast amount of training, both via the pre-training and the instruction-tuning stages, enables LLMs to perform complex tasks [10], perform in-context learning [26], etc. Recent advancements in LLMs have sparked simultaneous exploration and research into the applicability of these LLMs on a variety of different tasks, such as data labeling  [27, 28, 29], text classification  [30, 31], model explanation [7], etc. We add on to this emerging body of work and explore LLM-guided zero-shot counterfactual generation for NLP model evaluation.
LLMs for Counterfactual Generation
While the use of LLMs for generating counterfactuals is still an emerging direction, some recent works have started exploring the role and effectiveness of LLMs in counterfactual generation [32]. While  [13] provides a thorough evaluation on how prompting, model size, task complexity, etc. affect LLM generations of counterfactuals, authors in  [33] look at how large language models can be used to generate counterfactual data for training smaller language models. Authors in  [7] explore the use of LLMs for causal explainability using counterfactuals. Unlike prior work in this direction, in this paper we focus on exploring whether LLMs can be used to generate counterfactuals in the zero-shot setting, specifically to evaluate and stress test black-box model in a post-training, pre-deployment scenario.
III Zero-shot LLM-guided Counterfactual Generation
In this section, we describe our counterfactual generation methodology as shown in Figure 2. Following the causal explanation generation procedure in prior work  [7], we use state-of-the-art LLMs in an off-the-shelf manner, without any fine-tuning. We improve upon prior work  [7] by expanding and broadening their pipeline into a more general framework that can work for tasks other than causal explanation. Note that in this paper, we formulate and evaluate our pipeline on the broad task of text classification, whereas formulations for other text tasks can be derived in a similar manner. To facilitate this we explain the following components in our framework:
@@ III Zero-shot LLM-guided Counterfactual Generation
We leverage recent state-of-the-art LLMs as the counterfactual generators. Given that these models have been trained on vast amounts of textual data along with extensive instruction tuning, we assume that LLMs can learn to modify and perturb text input to simulate how human annotators generate counterfactuals for specific tasks  [34]. For this, we use both proprietary models from OpenAI and open-source models from Meta AI, and use carefully crafted instructions and constraint prompts to generate the counterfactuals. Specifically, we use the following proprietary models via the OpenAI API wherever applicable:
  • GPT-3.5333https://platform.openai.com/docs/models/gpt-3-5-turbo: Often referred to as ChatGPT. This is the model that has been explored in a variety of text applications. Specifically we use the gpt-3.5-turbo variant.
@@
  • GPT-4o [^5] https://platform.openai.com/docs/models/gpt-4o: This is the most recent and flagship model from the OpenAI GPT family of models. Although this model is capable of solving multimodal tasks (including text and visual input and text output), we only use it for text input. According to OpenAI, this model is capable of solving complex, multi-step tasks, thereby making it a suitable candidate for our task.
  • GPT-4o-mini [^6] https://platform.openai.com/docs/models/gpt-4o-mini: This is a low-cost, lower-latency and possibly smaller [^7] There is no official information regarding size of GPT-4o vs. GPT-4o-mini of the previous GPT-4o model. The lightweight nature of this model may be beneficial for researchers looking to use our pipeline for larger datasets.
Among the open-source models, we use the following models from Meta’s Llama family of models:
Due to resource constraints, we were unable to use larger variants of Llama 2 and Llama 3 models.
@@
To generate the counterfactuals in a zero-shot manner using the chosen LLM, the prompt needs to have informative instructions and constraints to guide the generation. We generate two types of counterfactuals: (1) actual counterfactuals: that is, counterfactuals that have a different label from the original input, according to the definition of counterfactual example. These are used in the counterfactual explanation experiments (see Section V), and (2) counterfactuals with same label as original: These are used in the contrast set experiments (see Section VI). For setting (1), we experiment with two variants of the generation process: (i) naive: Here the LLM is directly prompted to generate a counterfactual, and (ii) guided: Here we use a two-step process - first leveraging the LLM to identify the important input features (i.e., words) that result in the predicted label, and then prompting the same LLM to edit a minimal set of those identified features to generate the counterfactual. For ease of extraction of the generated counterfactuals, we also specify an output constraint that allows easy parsing based on a regular expression string match.
IV Experimental Settings
We use the pipeline described in the previous section in order to generate counterfactuals and demonstrate these for stress-testing and explaining black-box text classifiers. Specifically our tasks are: (1) Counterfactual explanations for explaining decisions of black-box text classifiers and (2) Evaluating black-box text classification models via contrast sets. To facilitate this, here we go over the datasets used and the general experimental setup for all our generation and evaluation experiments:
IV-A Datasets
@@
IV-B Experimental Setup
All experiments on open-source models were performed on two A100 GPUs with a total of 80G memory. For 13B Llama 2 models, we use 4-bit quantization using the optimal ‘nf4’ datatype [38]. For all LLMs, we use top_p sampling with $p=1$, temperature $t=0.4$ and a repetition penalty of $1.1$. We use PyTorch and fine-tuned models hosted on HuggingFace in both Sections V and VI.
V Counterfactual Explanations via LLM-generated Counterfactuals
Explainability is a major challenge in many NLP applications such as text classification [39, 40, 41]. Although recent models involving pre-trained transformer-based language models [42] have achieved or even exceeded human-level performance on several tasks [1, 2], most of these models are black-box by design and hence are not interpretable. Such models do not offer transparency on why it predicted a certain label, or even what features in the input resulted in the prediction. The ubiquity of these black-box classifiers necessitates the development of explanation frameworks and techniques that provide some degree of understanding into the decision-making function of the model [43]. Counterfactual explanations [4] give an insight into what could have been different in the input to change the output label predicted by the classifier. Gold-standard counterfactual generation requires human annotators and is also task-specific [44, 45], therefore making it an extremely expensive and labor-intensive endeavor. Therefore, we investigate whether we use zero-shot LLM-generated counterfactuals as counterfactual explanations for black-box text classifiers.